Q:-The non zero value of 'a' for which the lines 2x-y+3z+4=0=ax+y-z+2 and x-3y+z=0=x+2y+z+1 are coplanar is:


A. -2


B. 4


C. 6


D. 0

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

2x-y+3z+4=0=ax+y-z+2 and x-3y+z=0=x+2y+z+1 are coplanar

then point of intersection of line must satisfy all the 4 plane

solve plane

2x-y+3z+4=0

 x-3y+z=0

x+2y+z+1 =0

  x= 12/5

y = -1/5

z= -3

this point also satisfied by

ax-y+z+2=0

put value and find a

a=-2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin